OpenStudy (lgbasallote):

In electric fields..why is \(d\vec E = \frac{k\text d q}{r^2} (\hat r)\)

OpenStudy (anonymous):

you can often treat a charge distribution as a summation of infinitesimal point charges.

OpenStudy (anonymous):

so the field is the summation of their individual fields, by superposition.

OpenStudy (anonymous):

is that what you're asking?

OpenStudy (lgbasallote):

no. i meant where did that formula come from?

OpenStudy (anonymous):

Coulomb.

OpenStudy (lgbasallote):

isn't coulumb's law \[| \vec F | = \frac {kqq'}{r^2} (\hat r)\] how does that become dq?

OpenStudy (anonymous):

qE =F or F/q =E , whichever you prefer: if there's no charge at some location in a field, there's still a field. as for the 'dq', I went over that above.

OpenStudy (lgbasallote):

wait..where dod you get that?

OpenStudy (anonymous):

E, in this definition, is something like "the field that causes a force, F, on a charge q', located a distance r from the charge generating the field, q. "

OpenStudy (lgbasallote):

and q is the charge?

OpenStudy (anonymous):

yep.

OpenStudy (lgbasallote):

but that still doesn't explain why \[d\vec E = \frac {kdq}{r^2} (\hat r)\]

OpenStudy (anonymous):

does it explain why E = kq/r^2 ?

OpenStudy (anonymous):

because if it explains that, then it explains the differential form of that, since they're exactly the same but for scale.

OpenStudy (lgbasallote):

if i remember right.. \[\huge \vec E = \frac{\left ( \frac{kqq_o}{r^2} \right)}{q_o} \implies \frac{kq}{r^2}\] right?

OpenStudy (anonymous):

yep.

OpenStudy (lgbasallote):

now what?

OpenStudy (anonymous):

now make the charge tiny.

OpenStudy (anonymous):

the field is proportional to the charge.

OpenStudy (anonymous):

So you must make its field scale correspondingly.

OpenStudy (lgbasallote):

what do you mean make the charge tiny?

OpenStudy (lgbasallote):

do you mean take its derivative?

OpenStudy (anonymous):

x. dx. what's the difference?

OpenStudy (lgbasallote):

dx is the derivative of x?

OpenStudy (anonymous):

sort of. it's the infinitesimal of x. the differential form. if something is occurring on x, like say a Force, and everywhere on x that force is a bit different, then F(x)dx is the way you express it. F(x)x isn't going to usefully describe the situation.

OpenStudy (lgbasallote):

hmm so why only take the d of q and not others?

OpenStudy (anonymous):

k doesn't change. r you have to write as either constant or as a function of something. eg. for a line of charge along x, with a linear charge density of lambda, dq = lambda*dx and hopefully you can also express the distance (and direction) from each of those little 'dq's' to the point where you want to know the value and direction of the field in terms of x also...

OpenStudy (lgbasallote):

and q is just a variable?

OpenStudy (lgbasallote):

that's why it's the only one that's put with a d?

OpenStudy (anonymous):

it's a quantity.

OpenStudy (anonymous):

here:

OpenStudy (lgbasallote):

what do you mean by "a quantity"

OpenStudy (anonymous):